The weight of a species of cat follows the normal distribution with the mean of 32kg and the standard deviation
of 5kg.
1
Expert's answer
2020-12-08T09:54:24-0500

Initially the statement of the question was incomplete and we introduced additional details. We shall consider the following question: What is the probability that weight of a species of cat is at most 30 kg?


Answer:


"P(X\le30) = P(Z \le \frac {30 - \mu} {\sigma}) = P(Z \le \frac {30 - 32} {5}) = P(Z\le -0.4) = 0.3446"